Ariana took 2 hour longer to drive 360 miles on the first day of her trip then she took to drive 250 miles on the second day. If her speed was the same on both days, what was the driving time each day?
Answer by Crashy(2)

i am giving u the answer in unit of miles per hour.
first day =
data:
s = ??
t = 2 hours
distance = 360 miles
s = d/t
s = 360/2
s = 180 miles/hour
taking the speed same
second day =
DATA:
s = 180 miles/hour
distance = 250 miles
time = ??
s= d/t
t = s/d
putting the values
t = 180/250
t = 0.72 hours ANS
araina will take 0.72 hours to complete the travel of 250 miles on second day with same speed.